show that ( 1, 1, 1), ( 0, 1, 1) and ( 0, 1, -1) generate or spanned R^3
1
Expert's answer
2020-11-10T07:00:32-0500

Given Points are,

"(1,1,1),(0,1,1),(0,1,-1)" respectively.


Then "\Delta=\left|\begin{matrix}1&1&1\\0&1&1\\0&1&-1\\\end{matrix}\right|"


="[1(1(-1)-(1)(1))+0+0]\\1(-1-1)=-2"


Hence "\Delta\neq 0"

So Given point generate or spanned "R^3" .